# A Comprehensive Taxonomy of Negation for NLP and Neural Retrievers

### v1

@@ Abstract
Understanding and solving complex reasoning tasks is vital for addressing the information needs of a user. Although dense neural models learn contextualised embeddings, they still underperform on queries containing negation. To understand this phenomenon, we study negation in both traditional neural information retrieval and LLM-based models. We (1) introduce a taxonomy of negation that derives from philosophical, linguistic, and logical definitions; (2) generate two benchmark datasets that can be used to evaluate the performance of neural information retrieval models and to fine-tune models for a more robust performance on negation; and (3) propose a logic-based classification mechanism that can be used to analyze the performance of retrieval models on existing datasets. Our taxonomy produces a balanced data distribution over negation types, providing a better training setup that leads to faster convergence on the NevIR dataset. Moreover, we propose a classification schema that reveals the coverage of negation types in existing datasets, offering insights into the factors that might affect the generalization of fine-tuned models on negation.
@@ 1 Introduction
RQ1 aims to bring together research from the linguistic literature in a taxonomy on negation. We design our taxonomy to be exhaustive, with no overlap, relevant to IR tasks, and containing a balance of granularity that is appropriate for the AI community. To address RQ2, we propose two synthetically generated datasets that cover all proposed negation types. Figure 1 illustrates the task alongside the data type represented in our datasets. RQ3 analyzes the performance of neural IR models, providing insight into the gap between human understanding and LLM encoding of negation. RQ4 connects the taxonomy to formalizations that can be used as data classification mechanisms, which will allow us to study the distribution of negation in existing datasets and identify possible reasons why fine-tuning on such datasets does not guarantee a performance boost for neural IR models.
@@ 2 Motivation
Fine-tuning on negation datasets. One could argue that this problem can be mitigated through fine-tuning Dolci (2022). However, catastrophic forgetting occurs when a model is fine-tuned on a new dataset Hayes et al. (2019), even if its distribution is similar to the original training data. In certain cases, fine-tuning can lead to a degradation of performance in the original training set Peters et al. (2019); Merchant et al. (2020). Model sensitivity to parameter adjustments is particularly noticeable in information retrieval settings. This has been observed in traditional BERT-based architectures (Gerritse et al., 2022) and LLMs Soudani et al. (2024a). Although this behavior can be mitigated by freezing the model parameters and adding a language model head that is fine-tuned on a new dataset Huang et al. (2022); Lin et al. (2022), this method restricts the capabilities the model can learn. Weller et al. (2024) shows that fine-tuning on their proposed dataset (NevIR) leads to a noticeable decline in MSMarco generalization performance. This could be explained by catastrophic forgetting and fine-tuning on an overly specialized dataset.
Representations of negation. Another explanation for models under-performing on negation is the lack of a specialized pre-training setup, which can arise from an under-representation of negation in popular pre-training datasets Hossain et al. (2020), although LLMs have been trained on extensive crawled datasets and exhibit similar shortcomings. An improper training setup can also be caused by the training objective. While contrastive loss pushes representations that are different in content to be distant in the representation space, two negated statements are close in content while conveying opposite information. Hosseini et al. (2021); Noji and Takamura (2020) address the problem of having a misalignment between the training objective and the semantics of negation by proposing an ‘unlikelihood’ loss function used to further pre-train BERT on factually incorrect statements with negation cues.
@@ 4 Methodology
{forest}
for tree=
grow’=0,
rounded corners,
draw,
align=center,
text centered,
anchor=center,
baseline,
parent anchor=east,
child anchor=west,
edge=-latex, thick, draw=black!50,
font=,
fill=blue!5,
minimum height=0.7cm,
text width=2.4cm,
l sep=16mm,
s sep=3mm,
inner sep=2pt,
[Negation taxonomy, fill=gray!20, text width=3.2cm
[Logical operators, fill=blue!20
[Sentential
(no, not, none), draw=black, thick]
[Exclusion
[Exceptors
(other, besides), draw=black, thick] [Quantifiers, tier=ExclusionBottom [Contradiction
(for all..exists not), draw=black, thick]
[Contrary
(for all..not exists), draw=black, thick]
[Subcontradiction
(some.. some not), draw=black, thick]
[Affixal
(in-, il-), draw=black, thick]
[Lexical, fill=blue!20
[Implicit
(refuse, deny), draw=black, thick]
[Contrasting
[Immediate antonyms
(north, south), draw=black, thick]
[Mid antonyms
(fast, moderate), draw=black, thick]
[Polar antonyms
(fast, slow), draw=black, thick]
@@ 4.1 Taxonomy
Our primary classification criterion is on the scope (see Appendix A.1), distinguishing explicit negation realized by a logical operator Haegeman (1995), from lexical negation that is present through the semantics of the word itself Natayou (2014). Logical Operators append to a word or clause, reversing its meaning. In lexical negation, a word or phrase inherently evokes negation, without the need for an appended operator.
We identify three types of logical operators. Sentential Zeijlstra (2004) negation is signalled by sentential operators such as no, not and none, which have a fixed syntactic role and occupy defined positions within a sentence. Exclusion MacCartney and Manning (2008) is signalled by exclusionary operators that are either exceptors, such as besides and others (exceptors represent a unique type of negation, see more in Appendix A.2), or quantifiers, such as the universal quantifier for all and the existential quantifier exists. In Aristotelian logic Keenan and Westerståhl (1997); Horn (1989), these quantifiers define three fundamental relations: Contradiction, Contraries, and Subcontradiction. Finally, Affixal Zimmer (1966) negation is signalled by prefix and suffix operators that are preppended or appended to an existing word, such as: un-, in-, im-, il-, ir-, dis-, non-, mis-, ill-, -less, -free Wahyuni (2014).
@@ 4.2 Data Generation
We generate two synthetic datasets designed to cover all negation types described in the taxonomy. We construct the datasets as follows: (1) we prompt the LLM to generate 100 topics of general knowledge to ensure familiarity and avoid long-tail knowledge Askari et al. (2025); (2) for each topic, we ask the LLM to return one Wikipedia page that we check using the Wikipedia API, ensuring the generations are grounded in documented and factual information; (3) conditioned on a Wikipedia page, the LLM generates pairs and following the template of CondaQA Ravichander et al. (2022) and NevIR Weller et al. (2024). (3.1) Given detailed prompts constructed for the individual negation type, we ask the LLM to retrieve a paragraph that contains one specific negation as defined in the taxonomy. If the document does not contain explicit markers for the specified negation, the model will retrieve the closest match and rephrase it by injecting specific markers, i.e., keywords such as impossible instead of not possible. (3.2) Given the extracted paragraph, the LLM generates a query. This is the process of generating one pair . (3.3) For generating the second pair, we employ two strategies to produce different degrees of lexical overlap between the negated datasets. (1) Free Generation: generate a positive query by removing the negation from ; generate a positive document by answering . (2) Controlled Generation: generate a positive query by removing the negation from ; generate a positive document by removing the negation from . The two synthetically generated datasets have and instances, respectively, where a single instance has pairs and . Appendix A.3 provides the prompts used for generation, and an additional verification step for guaranteeing the relevance of documents; Table 3 and Figure 8 summarize the dataset statistics and distribution of generated labels.
@@ 4.3 LM Logic classification
Negation can be analysed at two granularities. Sentence-level: some negation types can be identified at the sentence level; if two sentences are either both negative or both positive, the pair agrees in polarity Mahany et al. (2022), and if they do not, it conveys a negative polarity relationship (sentential, exclusionary, affixal, and implicit). Pair-level: the negation polarity can only be identified by comparison, i.e., whether both statements can be true at the same time (quantifiers and contrasting negation). We propose a classification mechanism that assigns each instance in an existing dataset a category outlined in our taxonomy by converting it to natural logic using typed lambda calculus formalisations Barendregt (1985) (see Appendix A.2). We generate formalisations for each instance by prompting a model with an instruction to generate the typed lambda calculus proof, and return the predicates, quantifiers and -typed formula, following the prompt illustrated in Appendix A.3. We categorize an existing dataset in four iterative steps:
@@ 5 Experimental Setup
Evaluation of the generation. We assess the quality of the generated datasets with human annotation on 5% of the generations, with two annotators evaluating each instance on: (1) relevance of documents to each query, (2) presence of negation, (3) naturalness, (4) coherence, and (5) consistency of information within the document. The annotation was conducted with LabelStudio.111https://labelstud.io/ We assess the annotations on quantitative and qualitative measures, together with the annotator agreement. Appendix A.6 illustrates the questions for the annotators alongside further details for the setup. Tables 4 and 5 report the annotation metrics. The main findings are as follows:
Annotators reported 71-77 accuracy for document relevance and f1 score for negation presence.
On a scale of 1-5, the annotators reported an approximate quality of 4 on naturalness, coherence, and consistency of language.
@@
Retrieval Models. We study the performance of lexical, bi-encoder, crossencoder, late interaction and tranformer models trained for first-stage retrieval, ranking, sentence similarity, natural language inference (NLI) and next token prediction (NTP). We follow the experimental setup introduced by Elsen et al. (2025). We show the specifications of all models in Table 6 (Appendix A.7).
Metrics. The metric used to evaluate the task presented in Figure 1 is pairwise accuracy: for each instance queries and documents , the model independently ranks . The prediction is correct only when the system places above for and inverts the order for . Random performance for pairwise accuracy is .
Fine-tuning. We fine-tune three models: ColBERTv1, multi-qa-mpnet-base-dot-v1, and Mistral-7B-Instruct for 20 epochs on the free generated dataset and evaluate on NevIR Weller et al. (2024) and MSMarco Bajaj et al. (2016) dev data.
@@ 6.1 Evaluation on Synthetic Data
Figure 3 illustrates 20 models evaluated on the free generation dataset. Sparse, dual, and biencoders exhibit poor performance on all types of negation, except Sentence-T5: a dual encoder trained for semantic similarity. Both late-interaction and all cross-encoder models, except nli-deberta-v3-base, show strong performance on all negation types. BERT and T5-based cross-encoders perform better than models with a RoBERTa, ELECTRA, and DeBERTa backbone. All transformer-based models, except for Qwen 1.5B (which has a disadvantage in size, and it has been trained for NTP) perform well on almost all negation types.
@@ 6.2 Evaluation on Logic Filtered NevIR
When we apply the classification mechanism on the validation set of NevIR, we find that three main types of negation are present. Out of 225 pairs, correspond to , while have been classified as not containing negation which we label as Others, while the remaining pairs are spread across the other types of negation present in the taxonomy. These results are in line with hypothesis H3, which states that existing datasets have an uneven distribution of negation types.
@@ 6.3 Evaluation on Logic Filtered ExcluIR
When applying the classification mechanism on the ExcluIR test set, we find three types of negation: with pairs out of . Moreover, have been classified as “Other“ while are distributed among the other classes. This means that more than of the entire dataset has been classified as exclusionary. These results further support hypothesis H3.
As shown in Figure 12 (Appendix A.8), the performance of the model is approximately uniform between the three identified types of negation. This finding contradicts with out synthetic data experiments, where exclusionary negation was significantly more difficult to encode than the other types of negation. To further inspect the source of this discrepancy, we take a closer inspection of the ExcluIR instances identified as “Sentential” or “Implicit”. This reveals that these instances only have a different rephrasing of a task that essentially is still exclusion. One example extracted from the dataset is Can you tell me about Paul Ziert’s involvement in founding the Bart Conner Gymnastics Academy in Norman, Oklahoma, while avoiding any mention of Bart Conner’s role in the academy?. Our categorization mechanism identifies this instance as “Implicit”, while it has the form of a set subtraction, as per the definition of exceptors.
@@ 6.4 Fine-tuning
Evaluation partitions: We evaluate against the dev partition of NevIR that has 450 triplets (2 triplets = 1 pair).
@@
NevIR | .17 | .21 | .43 | .37 | .37 | .34 |
| Synth | .20 | .34 | .32 | .36 | .34 | .31 | |
| Mixed | .21 | .39 | .44 | .37 | .33 | .31 | |
@@
NevIR | .16 | .54 | .48 | .35 | .17 | .06 |
| Synth | .34 | .45 | .46 | .33 | .07 | .03 | |
| Mixed | .38 | .54 | .55 | .26 | .03 | .01 | |
@@
NevIR | .67 | .71 | .72 | .53 | .58 | .60 |
| Synth | .57 | .54 | .50 | .59 | .55 | .55 | |
| Mixed | .71 | .71 | .72 | .57 | .60 | .54 | |

### v2

@@ Abstract
Understanding and solving complex reasoning tasks is vital for addressing the information needs of a user. Although dense neural models learn contextualised embeddings, they underperform on queries containing negation. To understand this phenomenon, we study negation in traditional neural information retrieval and LLM-based models. We (1) introduce a taxonomy of negation that derives from philosophical, linguistic, and logical definitions; (2) generate two benchmark datasets that can be used to evaluate the performance of neural information retrieval models and to fine-tune models for a more robust performance on negation; and (3) propose a logic-based classification mechanism that can be used to analyze the performance of retrieval models on existing datasets. Our taxonomy produces a balanced data distribution over negation types, providing a better training setup that leads to faster convergence on the NevIR dataset. Moreover, we propose a classification schema that reveals the coverage of negation types in existing datasets, offering insights into the factors that might affect the generalization of fine-tuned models on negation. Our code is publicly available on GitHub111github.com/RoxanaPetcu/taxonomy-negation, and the datasets are available on HuggingFace222gpt4o-negation-controlled333gpt4o-negation-free.
@@ 1 Introduction
RQ1 aims to bring together research from the linguistic literature in a taxonomy on negation. We design our taxonomy to be exhaustive, with no overlap, and relevant to IR tasks. To address RQ2, we propose two synthetically generated datasets that cover all proposed negation types. Figure 1 illustrates the task alongside the data type represented in our datasets. RQ3 analyzes the performance of neural IR models, providing insight into the gap between human understanding and LLM encoding of negation. RQ4 connects the taxonomy to formalizations that can be used as data classification mechanisms, allowing to study existing datasets and identify reasons why fine-tuning does not guarantee a performance boost.
@@ 2 Motivation
Fine-tuning on negation datasets. One could argue that this problem can be mitigated through fine-tuning Dolci (2022). However, catastrophic forgetting occurs when a model is fine-tuned on a new dataset Hayes et al. (2019), even if its distribution is similar to the original training data. In certain cases, fine-tuning can lead to a degradation of performance in the original training set Peters et al. (2019); Merchant et al. (2020). Model sensitivity to parameter adjustments is particularly noticeable in information retrieval settings. This has been observed in traditional BERT-based architectures (Gerritse et al., 2022) and LLMs Soudani et al. (2024a). Although this behavior can be mitigated by freezing the model parameters and adding a language model head that is fine-tuned on a new dataset Huang et al. (2022); Lin et al. (2022), this method restricts the capabilities the model can learn. Weller et al. (2024) shows that fine-tuning on their proposed dataset (NevIR) leads to a noticeable decline in MSMarco generalization performance.
Representations of negation. Another explanation for models under-performing on negation is a lack of specialized pre-training setups, which can arise from an under-representation of negation in popular pre-training datasets Hossain et al. (2020), although LLMs have been trained on extensive crawled datasets and exhibit similar shortcomings. An improper training can also be caused by the training objective. While contrastive loss pushes different content to be distant in the representation space, two negated statements are close in content while conveying opposite information. Hosseini et al. (2021); Noji and Takamura (2020) address the problem of having a misalignment between the training objective and the semantics of negation by proposing an ‘unlikelihood’ loss function used to further pre-train BERT on factually incorrect statements with negation cues.
@@ 4 Methodology
[…]
@@ 4.1 Taxonomy
Our primary classification criterion is on the scope of negation (the part of a sentence whose meaning is altered by negation), distinguishing explicit negation realized by a logical operator Haegeman (1995), from lexical negation that is present through the semantics of the word itself Natayou (2014). Logical Operators append to a word or clause, reversing its meaning. In lexical negation, a word or phrase inherently evokes negation, without the need for an appended operator.
We identify three types of logical operators based on literature review Horn (1989). Sentential Zeijlstra (2004) negation is signalled by sentential operators such as no, not and none, which have a fixed syntactic role and occupy defined positions within a sentence. Exclusion MacCartney and Manning (2008) is signalled by exclusionary operators that are either exceptors, such as besides and others (exceptors represent a unique type of negation, see more in Appendix A.2), or quantifiers, such as the universal quantifier for all and the existential quantifier exists. In Aristotelian logic Keenan and Westerståhl (1997); Horn (1989), these quantifiers define three fundamental relations: Contradiction, Contraries, and Subcontradiction. Finally, Affixal Zimmer (1966) negation is signalled by prefix and suffix operators that are preppended or appended to an existing word, such as: un-, in-, im-, il-, ir-, dis-, non-, mis-, ill-, -less, -free Wahyuni (2014).
@@ 4.2 Data Generation
We generate two synthetic datasets designed to cover all negation types described in the taxonomy. We construct the datasets as follows: (1) we prompt an LLM to generate 100 topics of general knowledge to ensure familiarity Askari et al. (2025) and avoid long-tail knowledge; (2) for each topic, we ask the LLM to return one Wikipedia page that we check using the Wikipedia API, ensuring the generations are grounded in documented and factual information; (3) conditioned on a Wikipedia page, the LLM generates pairs and following the template of CondaQA Ravichander et al. (2022) and NevIR Weller et al. (2024). (3.1) Given detailed prompts constructed for the individual negation type, we ask the LLM to retrieve a paragraph that contains one specific negation as defined in the taxonomy. If the document does not contain explicit markers for the specified negation, the model will retrieve the closest match and rephrase it by injecting specific markers, i.e., keywords such as impossible instead of not possible. This phenomenon was observed with affixal negations, which our approach translated as a sentential one, as they are guaranteed to be semantically equivalent. The other types of negation that were not always present in the documents were the quantifiers, which can be translated from one to the other with logic transformations. (3.2) Given the extracted paragraph, the LLM generates a query. This is the process of generating one pair . (3.3) For generating the second pair, we employ two strategies to produce different degrees of lexical overlap between the negated datasets. (1) Free Generation: generate a positive query by removing the negation from ; generate a positive document by answering . (2) Controlled Generation: generate a positive query by removing the negation from ; generate a positive document by removing the negation from . The two synthetically generated datasets have and instances, respectively, where a single instance has pairs and . Appendix A.3 provides the prompts used for generation, and an additional verification step for guaranteeing the relevance of documents; Table 4 and Figure 8 summarize the dataset statistics and distribution of generated labels.
@@ 4.3 LM Logic classification
Negation can be analysed at two granularities. Sentence-level: some negation types can be identified at the sentence level; if two sentences are either both negative or both positive, the pair agrees in polarity Mahany et al. (2022), and if they do not, it conveys a negative polarity relationship (sentential, exclusionary, affixal, and implicit). Pair-level: the negation polarity can only be identified by comparison, i.e., whether both statements can be true at the same time (quantifiers and contrasting negation). We propose a classification mechanism that assigns each instance in an existing dataset a category outlined in our taxonomy by converting it to natural logic using typed lambda calculus formalisations Barendregt (1985) (see Appendix A.2). We generate formalisations for each instance by prompting a model with an instruction to generate the typed lambda calculus proof, and return the predicates, quantifiers and -typed formula. We categorize an existing dataset in four iterative steps:
@@ 5 Experimental Setup
| BM25 | Sparse | Retrieval | N/A | N/A | N/A |
| DPR [29] | Bi-Encoder | Retrieval | NQ | 219M | BERT |
| coCondenser [14] | Bi-Encoder | Retrieval | MSMarco | 110M | BERT |
| Dragon [36] | Bi-Encoder | Retrieval | MS MARCO | N/A | BERT |
| Sentence-T5 | Dual Encoder | Sentence Similarity | NLI | 220M | T5 |
| ColBERTv1 [32] | Late Interaction | Retrieval | MSMarco | 110M | BERT |
| ColBERTv2 [58] | Late Interaction | Retrieval | MSMarco | 110M | BERT |
| MonoT5 Base [51] | Crossencoder | Ranking | MSMarco | 223M | T5 |
| MonoT5 Large [51] | Crossencoder | Ranking | MSMarco | 737M | T5 |
| MonoT5 3B [51] | Crossencoder | Ranking | MSMarco | 2.85B | T5 |
| qnli-electra-base | Crossencoder | NLI | QNLI | 110M | ELECTRA |
| Mistral-7B-Instruct [26] | Transformer | NTP | Crawled | 7B | BPE |
| Llama-3.1-3B-Instruct [16] | Transformer | NTP | Crawled | 7B | Llama |
| Llama-3.2-8B-Instruct [16] | Transformer | NTP | Crawled | 7B | Llama |
Evaluation of the generation. We assess the quality of the generated datasets with human annotation on 5% of the generations, with two annotators evaluating each instance on: (1) relevance of documents to each query, (2) presence of negation, (3) naturalness, (4) coherence, and (5) consistency of information within the document. The annotation was conducted with LabelStudio.444https://labelstud.io/ We assess the annotations on quantitative and qualitative measures, together with the annotator agreement. Appendix A.6 illustrates the questions for the annotators, metrics used, alongside further details for the setup. For both performance and inner annotator agreement, we use metrics such as f1-score, average on ordinal scales, and (weighted) Cohen’s Kappa. Tables 5 and 6 report the annotation metrics. The main findings are as follows:
Annotators reported 71–77 accuracy for document relevance and 83%–88% f1 score for negation presence.
On a scale of 1–5, the annotators reported an approximate quality of 4 on naturalness, coherence, and consistency of language.
@@
Human performance on the synthetic datasets shows a pairwise accuracy score of for free generation, and for controlled generation on identifying the relevant document for each question.
@@
Retrieval Models. We study the performance of lexical, bi-encoder, cross-encoder, late interaction and transformer models trained for first-stage retrieval, ranking, sentence similarity, natural language inference (NLI) and next token prediction (NTP). We follow the experimental setup introduced by Elsen et al. (2025). We show the specifications of all models in Table 1.
Datasets. We evaluate on three benchmarks. NevIR and ExcluIR are two contrastive benchmarks where each instance comprises of two documents and two queries that only differ by a targeted negation, or exclusion. We also use MSMarco dev partition, which is not specifically designed for contrastive pairs, but is used simply as a complex retrieval benchmark.
Metrics. The metric used to evaluate the task is pairwise accuracy: for each instance queries and documents , the model independently ranks . The prediction is correct only when the system places above for and inverts the order for . Random performance for pairwise accuracy is .
Fine-tuning. We fine-tune three models: ColBERTv1, multi-qa-mpnet-base-dot-v1, and Mistral-7B-Instruct for 20 epochs on the free generated dataset and evaluate on NevIR Weller et al. (2024) test and MSMarco Bajaj et al. (2016) dev data.
@@ 6.1 Evaluation on Synthetic Data
Figure 3 illustrates 20 models evaluated on the free generation dataset. Sparse, dual, and biencoders exhibit poor performance on all types of negation, except Sentence-T5: a dual encoder trained for semantic similarity. Both late-interaction and all cross-encoder models, except nli-deberta-v3-base, show strong performance on all negation types. BERT and T5-based cross-encoders perform better than models with a RoBERTa, ELECTRA, and DeBERTa backbone. All transformer-based models, except for Qwen 1.5B (which has a disadvantage in size, and which has been trained for NTP) perform well on almost all negation types.
@@ 6.2 Evaluation on Logic Filtered NevIR
When we apply the classification mechanism on the validation set of NevIR, we find that three main types of negation are present. Out of 225 pairs, correspond to , while have been classified as not containing negation, in which case we label as Others, while the remaining pairs are spread across the other types of negation present in the taxonomy. These results are in line with hypothesis H3, which states that existing datasets have an uneven distribution of negation types.
@@ 6.3 Evaluation on Logic Filtered ExcluIR
When applying the classification mechanism on the ExcluIR test set, we find three types of negation: with pairs out of . Moreover, have been classified as “Other” while are distributed among the other classes. This means that more than of the entire dataset has been classified as exclusionary. These results further support hypothesis H3.
As shown in Figure 12 (Appendix A.7), the performance of the model is approximately uniform between the three identified types of negation. This finding contradicts with our synthetic data experiments, where exclusionary negation was significantly more difficult to encode than the other types of negation. To further inspect the source of this discrepancy, we take a closer inspection of the ExcluIR instances identified as “Sentential” or “Implicit”. This reveals that these instances only have a different rephrasing of a task that essentially is still exclusion. One example extracted from the dataset is ‘Can you tell me about Paul Ziert’s involvement in founding the Bart Conner Gymnastics Academy in Norman, Oklahoma, while avoiding any mention of Bart Conner’s role in the academy?’. Our categorization mechanism identifies this instance as “Implicit”, while it has the form of a set subtraction, as per the definition of exceptors.
@@ 6.4 Fine-tuning
Evaluation partitions: We evaluate against the test partition of NevIR that has 2.8k triplets (2 triplets = 1 pair), and against the dev partition of MSMarco.
@@
NevIR | .21 | .24 | .45 | .37 | .37 | .34 |
| Synth | .23 | .33 | .36 | .36 | .34 | .31 | |
| Mixed | .23 | .40 | .48 | .37 | .33 | .31 | |
@@
NevIR | .12 | .51 | .52 | .35 | .17 | .06 |
| Synth | .34 | .38 | .40 | .33 | .07 | .03 | |
| Mixed | .36 | .52 | .50 | .26 | .03 | .01 | |
@@
NevIR | .70 | .78 | .78 | .53 | .58 | .60 |
| Synth | .58 | .58 | .58 | .59 | .55 | .55 | |
| Mixed | .72 | .78 | .78 | .57 | .60 | .54 | |